# Breaking the Exploration Bottleneck: Rubric-Scaffolded Reinforcement Learning for General LLM Reasoning

### v5

@@ 5.1 Experimental Setups
Table 1: Main results comparison with closed-source and open-source models across different benchmarks. The best results in each box are highlighted in bold.
@@
Qwen3-30B-A3B-Instruct 46.8 71.4 84.2 71.3 78.4 74.4 83.0 31.9
+ RuscaRL 61.1 (+14.3) 73.0 (+1.6) 84.8 (+0.6) 71.9 (+0.6) 79.2 (+0.8) 74.3 (-0.1) 84.5 (+1.5) 32.1 (+0.2)
Qwen3-30B-A3B-Base 11.4 43.3 73.6 65.1 36.8 35.8 39.0 13.3
+ RuscaRL 48.3 (+36.9) 60.9 (+17.6) 71.3 (-2.3) 65.4 (+0.3) 59.0 (+22.2) 46.0 (+10.2) 76.3 (+37.3) 30.3 (+17.0)
Qwen2.5-7B-Instruct 23.6 47.9 61.8 56.3 45.3 37.4 71.0 26.8
+ RuscaRL 50.3 (+26.7) 61.2 (+13.3) 63.5 (+1.7) 56.5 (+0.2) 56.3 (+11.0) 38.6 (+1.2) 75.3 (+4.3) 31.0 (+4.2)
Qwen2.5-7B 8.2 28.2 55.3 55.0 23.5 30.3 32.0 14.5
+ RuscaRL 46.2 (+38.0) 47.8 (+19.6) 58.2 (+2.9) 55.6 (+0.6) 45.8 (+22.3) 34.8 (+4.5) 56.2 (+24.2) 25.9 (+11.4)
Llama-3.1-8B-Instruct 12.4 29.8 66.8 58.0 36.7 44.5 72.6 22.6
+ RuscaRL 46.0 (+33.6) 46.3 (+16.5) 70.7 (+3.9) 60.7 (+2.7) 52.7 (+16.0) 54.2 (+9.7) 79.7 (+7.1) 31.1 (+8.5)
Llama-3.1-8B 0 9.1 36.9 35.9 12.9 26.3 18.1 11.6
+ RuscaRL 25.8 (+25.8) 29.5 (+20.4) 49.7 (+12.8) 45.4 (+9.5) 35.6 (+22.7) 33.3 (+7.0) 55.6 (+37.5) 21.4 (+9.8)
Figure 3: STEM Benchmarks.
@@
Models and Training Settings. We use multiple initial models from different series and parameter scales for our experiments, including both instruct models and base models from the Qwen2.5 series (Yang et al., 2024), the Qwen3 series (Yang et al., 2025), and the Llama-3 series (Meta-AI, 2025; Grattafiori et al., 2024). All models are trained using the GRPO algorithm on the verl framework (Sheng et al., 2025). Detailed training settings are provided in Appendix C.1.
@@
Evaluation Benchmarks. We use HealthBench-500, a randomly selected subset of 500 samples from HealthBench (Arora et al., 2025), as our held-out evaluation set. Additionally, we evaluate on other medical benchmarks including LLMEval-Med (Zhang et al., 2025b), MedQA (Jin et al., 2021), and MedMCQA (Pal et al., 2022). For the Writing domain, we use WritingBench (Wu et al., 2025c) and Creative Writing v3 (Paech, 2025) benchmarks. For the Instruction Following domain, we use IFEVAL (Zhou et al., 2023b) and IFBench (Pyatkin et al., 2025) benchmarks. For the STEM domain, we use GPQA-Diamond (Rein et al., 2024), MMLU (Hendrycks et al., 2020), MMLU-Pro (Wang et al., 2024), MATH-500 (Lightman et al., 2023), AMC 2023111https://huggingface.co/datasets/knoveleng/AMC-23, AIME 2024, and AIME 2025222https://artofproblemsolving.com/wiki/index.php/AIME_Problems_and_Solutions. Detailed evaluation settings are provided in Appendix C.2.
@@
Baselines. We compare RuscaRL against four representative baseline methods: (1) Rubric-based RL: A rubric-as-reward RL baseline implemented with GRPO algorithm using rubrics scores as rewards (Gunjal et al., 2025). (2) Rubric-based RL with full scaffolding (Rubric-based RL-S): A method that provides all rubrics with scaffolding support in the instruction, without intra‑group differentiation and without an inter‑step decay function. (3) SFT: Fine-tuned on GPT-4.1 (OpenAI, 2025b) demonstrations generated with scaffolding support. (4) SFT + Rubric-based RL: A combination approach that first applies SFT and then applies rubric-based RL training.
@@ 5.2 Overall Performance
RuscaRL achieves consistent and notable gains across tasks and model scales, showcasing its effectiveness and broad generalization. Across medical, writing, and instruction-following tasks (Table 1), RuscaRL achieves significant gains over multiple initial models, with Qwen3-30B-A3B-Instruct on HealthBench-500 surpassing many leading models (e.g., OpenAI-o3). Notably, RuscaRL is particularly effective on instruct models and provides larger benefits for weaker models, such as Llama-3.1-8B-Instruct. This advantage stems from our scaffolding approach, which leverages the model’s existing instruction-following ability to elicit higher-quality and more diverse responses, explaining why RuscaRL is especially well-suited for training on models with strong instruction-following capacities. Meanwhile, RuscaRL has also been successfully extended to the STEM domain: experiments on Qwen2.5-7B-Instruct show consistent performance improvements across all STEM benchmarks (see Figure 3). More detailed results about performance across different model series and scales are provided in Appendix D.1, which further demonstrates the robustness and broad applicability of our approach. Additionally, we explore the effects of mixing training data from different domains in Appendix D.2.
Table 2: Performance comparison of different baseline methods and RuscaRL across multiple domains. Bold indicates the best performance of each method.
@@
Initial 23.6 47.9 61.8 56.3 45.3 37.4 71.0 26.8
Rubric-based RL 41.2 (+17.6) 54.6 (+6.7) 62.1 (+0.3) 56.3 (+0.0) 53.7 (+8.4) 38.8 (+1.4) 75.1 (+4.1) 29.3 (+2.5)
Rubric-based RL-S 36.6 (+13.0) 56.1 (+8.2) 57.9 (-3.9) 52.4 (-3.9) 45.8 (+0.5) 38.3 (+0.9) 71.9 (+0.9) 28.6 (+1.8)
RuscaRL (Ours) 50.3 (+26.7) 61.2 (+13.3) 63.5 (+1.7) 56.5 (+0.2) 56.3 (+11.0) 38.6 (+1.2) 75.3 (+4.3) 31.0 (+4.2)
SFT 38.3 (+14.7) 52.6 (+4.7) 60.8 (-1.0) 57.3 (+1.0) 62.8 (+17.5) 45.3 (+7.9) 75.2 (+4.2) 25.2 (-1.6)
SFT + Rubric-based RL 55.1 (+31.5) 58.5 (+10.6) 59.7 (-2.1) 56.4 (+0.1) 66.7 (+21.4) 43.6 (+6.2) 82.1 (+11.1) 29.6 (+2.8)
SFT + RuscaRL (Ours) 56.9 (+33.3) 58.8 (+10.9) 61.6 (-0.2) 56.9 (+0.6) 66.8 (+21.5) 43.9 (+6.5) 82.5 (+11.5) 30.6 (+3.8)
@@
Initial 8.2 28.2 55.3 55.0 23.5 30.3 32.0 14.5
Rubric-based RL 41.9 (+33.7) 46.5 (+18.3) 48.2 (-7.1) 49.9 (-5.1) 40.2 (+16.7) 33.8 (+3.5) 53.4 (+21.4) 25.5 (+11.0)
Rubric-based RL-S 21.7 (+13.5) 44.4 (+16.2) 60.3 (+5.0) 55.5 (+0.5) 43.9 (+20.4) 25.7 (-4.6) 52.3 (+20.3) 20.4 (+5.9)
RuscaRL (Ours) 46.2 (+38.0) 47.8 (+19.6) 58.2 (+2.9) 55.6 (+0.6) 45.8 (+22.3) 34.8 (+4.5) 56.2 (+24.2) 25.9 (+11.4)
SFT 32.2 (+24.0) 40.0 (+11.8) 56.5 (+1.2) 54.4 (-0.6) 56.6 (+33.1) 42.5 (+12.2) 69.7 (+37.7) 20.8 (+6.3)
SFT + Rubric-based RL 36.4 (+28.2) 39.7 (+11.5) 57.1 (+1.8) 54.1 (-0.9) 57.1 (+33.6) 43.2 (+12.9) 71.6 (+39.6) 23.7 (+9.2)
SFT + RuscaRL (Ours) 35.4 (+27.2) 42.7 (+14.5) 58.2 (+2.9) 55.1 (+0.1) 57.4 (+33.9) 42.6 (+12.3) 72.0 (+40.0) 23.1 (+8.6)
RuscaRL consistently outperforms Rubrics-based methods across tasks. As shown in Table 2, in the direct RL setting, RuscaRL achieves the best performance on most medical, writing, and instruction-following tasks, delivering larger and more stable gains than Rubric-based RL and RL-S (e.g., 50.3 vs. 41.2 and 36.6 accuracy on HealthBench-500 with Qwen2.5-7B-Instruct). In the SFT-then-RL setting, both RuscaRL and Rubric-based RL achieve additional improvements over SFT, with RuscaRL generally yielding larger gains across most tasks, though the magnitude is smaller than in the direct RL setting. We argue that RuscaRL essentially leverages rubrics as prior knowledge to guide exploration, while SFT also serves to accelerate RL exploration (cold start). Since both mechanisms overlap in facilitating exploration, this may explain why the performance gap between RuscaRL and Rubric-based RL narrows under the SFT-then-RL setting.
@@ 5.3 Analysis
In this subsection, we conduct an analysis of RuscaRL using Qwen2.5-7B-Instruct as the initial model and HealthBench as the training and evaluation dataset. In addition, we compare three approaches in the following analysis: RuscaRL, RuscaRL (RuscaRL without the inter-step decay mechanism), and Rubric-based RL. We use the Best-of-N [^3] For cost considerations in the Best-of-N evaluation, we use Qwen3-32B (non-thinking) as the Grader LLM. metric to reflect both the model’s reasoning boundaries (at large N) and sampling efficiency (at small N).
Figure 4: Best-of-N Performance.
RuscaRL significantly improves sampling efficiency and reasoning boundaries. As shown in Figure 4, RuscaRL significantly improves single-sample quality at N=1, indicating that the scaffolding mechanism effectively enhances the model’s reasoning stability. At N=2048, its performance ceiling surpasses both the initial model and Rubric-based RL, validating its advantage in expanding the reasoning boundary. Moreover, RuscaRL exhibits a steeper performance curve across N, meaning it can achieve the same performance with fewer samples. Further analysis (Appendix D.4.1) indicates that RuscaRL also produces highly novel responses that the initial model could barely generate, showing that rubric scaffolding effectively breaks the exploration bottleneck and discovers new solutions.
RuscaRL achieves exploration-exploitation balance. As shown in Figure 5(a), RuscaRL demonstrates a well-balanced exploration–exploitation trajectory: policy entropy first rises as the model explores diverse reasoning trajectories, then declines as it converges to high-quality patterns. In contrast, RuscaRL suffers from uncontrolled entropy growth leading to instability, while Rubric-based RL collapses under continuous entropy decline. Validation accuracy (Figure 5(b)) consistently shows RuscaRL achieving the best performance throughout training, demonstrating long-term stability without policy entropy collapses, followed by Rubric-based RL and then RuscaRL. Similar trends are observed in Self-BLEU and Semantic Distance (Appendix D.4.2), confirming that RuscaRL achieves effective exploration followed by stable exploitation.
(a) Training policy entropy
(b) Validation evaluation score
(c) Sigmoid decay function
Figure 5: Training dynamics. The figure shows the evolution of policy entropy, validation accuracy, and sigmoid decay function during training.
@@ 5.4 Ablation Studies
Inter-step Decay Analysis. We analyze different decay functions for inter-step control during training. We define the base scaffolding intensity of inter-step control as $f(t)$, where $t$ is the normalized training progress ($t\in[0,1]$). We compare the following decay functions: (1) Sigmoid (Ours): S-shaped decay function $f(t)=\frac{1}{1+e^{\alpha(t-t_{0})}}$, where parameter $\alpha$ controls the steepness of decay and $t_{0}$ controls the midpoint of decay, achieving smooth nonlinear transitions. (2) Constant: Constant control $f(t)=1$, maintaining constant full scaffolding. (3) Linear: Linear decay function f(t)=1tf(t)=1-t, achieving uniform linear decrease. (4) Power (n): Power decay function $f(t)=(1-t)^{n}$, where n controls the curvature of decay, including various power configurations.
@@
Based on the superior performance of the sigmoid function, we further analyze the effects of both parameter dimensions (speed $\alpha$ and midpoint $t_{0}$) using Qwen2.5-7B-Instruct as the initial model and HealthBench as the training and evaluation dataset. Figures 6(c) and 6(d) demonstrate the performance differences across various sigmoid parameter configurations, ultimately determining the optimal configuration as $\alpha=125,t_{0}=0.2$. (1) Removing scaffolding too fast (large $\alpha$) prevents the model from adapting to rapid scaffolding changes, easily causing training instability; while removing scaffolding too slowly (small $\alpha$) leads to incomplete early-stage scaffolding, failing to fully stimulate the model’s exploration capability, and prolonged retention of scaffolding in later stages also causes overfitting issues. (2) Starting decay too early (small $t_{0}$) leads to insufficient scaffolding support, causing the model to lack necessary guidance in early training stages; while starting decay too late (large $t_{0}$) causes the model to over-rely on scaffolding, ultimately resulting in overfitting.

### v6

@@ 5.1 Experimental Setups
Table 1: Main results comparison with baseline models across different benchmarks. The best results in each box are highlighted in bold.
@@
Qwen3-30B-A3B-Instruct 46.9 71.5 84.2 71.3 78.1 74.4 83.0 31.9
+ RuscaRL 61.1 (+14.2) 73.2 (+1.7) 84.8 (+0.6) 71.9 (+0.6) 79.2 (+1.1) 74.3 (-0.1) 84.5 (+1.5) 32.1 (+0.2)
Qwen3-30B-A3B-Base 11.2 43.1 73.6 65.1 36.9 35.8 39.0 13.3
+ RuscaRL 48.4 (+37.2) 60.9 (+17.8) 71.3 (-2.3) 65.4 (+0.3) 59.5 (+22.6) 46.0 (+10.2) 76.3 (+37.3) 30.3 (+17.0)
Qwen2.5-7B-Instruct 23.4 48.0 61.8 56.3 45.2 37.4 71.0 26.8
+ RuscaRL 56.4 (+33.0) 65.3 (+17.3) 63.5 (+1.7) 56.5 (+0.2) 56.1 (+10.9) 38.6 (+1.2) 75.3 (+4.3) 31.0 (+4.2)
Qwen2.5-7B 8.5 28.2 55.3 55.0 23.8 30.3 32.0 14.5
+ RuscaRL 46.3 (+37.8) 47.9 (+19.7) 58.2 (+2.9) 55.6 (+0.6) 46.0 (+22.2) 34.8 (+4.5) 56.2 (+24.2) 25.9 (+11.4)
Llama-3.1-8B-Instruct 12.5 30.1 66.8 58.0 36.7 44.5 72.6 22.6
+ RuscaRL 46.0 (+33.5) 46.2 (+16.1) 70.7 (+3.9) 60.7 (+2.7) 52.7 (+16.0) 54.2 (+9.7) 79.7 (+7.1) 31.1 (+8.5)
Llama-3.1-8B 0.0 9.1 36.9 35.9 13.0 26.3 18.1 11.6
+ RuscaRL 25.8 (+25.8) 29.6 (+20.5) 49.7 (+12.8) 45.4 (+9.5) 35.7 (+22.7) 33.3 (+7.0) 55.6 (+37.5) 21.4 (+9.8)
@@
Models and Training Settings. We use several backbone models from different series and with varying parameter scales for our experiments, including both instruction-tuned models and base models from the Qwen2.5 series (Yang et al., 2024), the Qwen3 series (Yang et al., 2025), and the Llama-3 series (Dubey et al., 2024). All these models are trained using the GRPO algorithm on the verl framework (Sheng et al., 2025). Detailed training settings are provided in Appendix C.1.
@@
Evaluation Benchmarks. We use HealthBench-500, a randomly selected subset of 500 samples from HealthBench (Arora et al., 2025), as our held-out evaluation set. Additionally, we evaluate on other medical benchmarks including LLMEval-Med (Zhang et al., 2025b), MedQA (Jin et al., 2021), and MedMCQA (Pal et al., 2022). For the Writing domain, we use WritingBench (Wu et al., 2025c) and Creative Writing v3 (Paech, 2025) benchmarks. For the Instruction Following domain, we use IFEVAL (Zhou et al., 2023c) and IFBench (Pyatkin et al., 2025) benchmarks.
@@
Baselines. We compare RuscaRL against several representative baseline methods: (1) Rubric-based RL: A rubric-as-reward RL ablation baseline implemented with the GRPO algorithm using rubric scores as rewards, corresponding to Rubric as Verifiable Rewards for Exploitation without scaffolding mechanisms. (2) MeRF (Zhang et al., 2025a): Following the MeRF paradigm of injecting reward specifications as in-context motivation, we include all rubrics in the prompt as scaffolding throughout training, without intra-group differentiation or inter-step decay. (3) RL-Plus (Dong et al., 2025): A hybrid-policy RL method that combats capability boundary collapse through mixture-policy importance sampling and exploration-oriented advantage shaping, encouraging continued discovery of correct but underrepresented reasoning paths. (4) Entropy-Augmented RL (EA-RL) (Cheng et al., 2025): An advantage-shaping approach that adds a stabilized entropy bonus to encourage exploratory reasoning and mitigate exploration collapse. (5) SFT: Fine-tuned on GPT-4.1 (OpenAI, 2025a) demonstrations generated with scaffolding support. (6) SFT + Rubric-based RL: A combination approach that first applies SFT and then applies rubric-based RL training.
@@ 5.2 Overall Performance
RuscaRL achieves consistent and notable gains across tasks and model scales, showcasing its effectiveness and broad generalization. Across medical, writing, and instruction-following tasks (Table 1), RuscaRL achieves significant gains over multiple initial models, with Qwen3-30B-A3B-Instruct on HealthBench-500 surpassing many leading models (e.g., OpenAI-o3). It is worth noting that our training data consists of open-ended tasks, whereas MedQA and MedMCQA are closed-ended multiple-choice benchmarks. Our improvements on these closed-ended benchmarks are marginal and are included only to demonstrate cross-task generalization. Notably, RuscaRL is particularly effective on instruction-tuned models and provides larger benefits for weaker models, such as Llama-3.1-8B-Instruct. This advantage stems from our scaffolding approach, which leverages the model’s existing instruction-following ability to elicit higher-quality and more diverse responses, explaining why RuscaRL is especially well-suited for training on models with strong instruction-following capacities. More detailed results about performance across different model series and scales are provided in Appendix D.1, which further demonstrates the robustness and broad applicability of our approach. Additionally, we explore the effects of mixing training data from different domains in Appendix D.3.
Table 2: Performance comparison of baseline methods across different benchmarks. The best results in each box are highlighted in bold.
@@
Initial 23.4 48.0 61.8 56.3 45.2 37.4 71.0 26.8
Rubric-based RL 52.0 (+28.6) 56.5 (+8.5) 62.3 (+0.5) 56.3 (+0.0) 53.7 (+8.5) 38.8 (+1.4) 75.1 (+4.1) 29.3 (+2.5)
MeRF 36.8 (+13.4) 56.1 (+8.1) 57.9 (-3.9) 52.4 (-3.9) 45.9 (+0.7) 38.3 (+0.9) 71.9 (+0.9) 28.6 (+1.8)
RL-Plus 53.6 (+30.2) 58.2 (+10.2) 62.7 (+0.9) 56.4 (+0.1) 54.8 (+9.6) 38.5 (+1.1) 75.0 (+4.0) 29.8 (+3.0)
EA-RL 51.3 (+27.9) 57.8 (+9.8) 61.5 (-0.3) 56.5 (+0.2) 52.9 (+7.7) 39.2 (+1.8) 74.6 (+3.6) 28.7 (+1.9)
RuscaRL (Ours) 56.4 (+33.0) 65.3 (+17.3) 63.5 (+1.7) 56.5 (+0.2) 56.1 (+10.9) 38.6 (+1.2) 75.3 (+4.3) 31.0 (+4.2)
SFT 38.3 (+14.9) 52.6 (+4.6) 60.8 (-1.0) 57.3 (+1.0) 62.8 (+17.6) 45.3 (+7.9) 75.2 (+4.2) 25.2 (-1.6)
SFT + Rubric-based RL 55.5 (+32.1) 58.5 (+10.5) 59.7 (-2.1) 56.4 (+0.1) 66.7 (+21.5) 43.6 (+6.2) 82.1 (+11.1) 29.6 (+2.8)
SFT + RuscaRL (Ours) 56.9 (+33.5) 58.8 (+10.8) 61.6 (-0.2) 56.9 (+0.6) 67.0 (+21.8) 43.9 (+6.5) 82.5 (+11.5) 30.6 (+3.8)
@@
Initial 8.5 28.2 55.3 55.0 23.8 30.3 32.0 14.5
Rubric-based RL 42.0 (+33.5) 46.5 (+18.3) 48.2 (-7.1) 49.9 (-5.1) 40.1 (+16.3) 33.8 (+3.5) 53.4 (+21.4) 25.5 (+11.0)
MeRF 21.7 (+13.2) 44.4 (+16.2) 60.3 (+5.0) 55.5 (+0.5) 43.4 (+19.6) 25.7 (-4.6) 52.3 (+20.3) 20.4 (+5.9)
RL-Plus 43.8 (+35.3) 47.2 (+19.0) 49.6 (-5.7) 50.8 (-4.2) 42.3 (+18.5) 34.2 (+3.9) 54.6 (+22.6) 25.7 (+11.2)
EA-RL 41.2 (+32.7) 45.8 (+17.6) 47.5 (-7.8) 51.2 (-3.8) 41.5 (+17.7) 33.2 (+2.9) 54.1 (+22.1) 24.8 (+10.3)
RuscaRL (Ours) 46.3 (+37.8) 47.9 (+19.7) 58.2 (+2.9) 55.6 (+0.6) 46.0 (+22.2) 34.8 (+4.5) 56.2 (+24.2) 25.9 (+11.4)
SFT 32.2 (+23.7) 40.0 (+11.8) 56.5 (+1.2) 54.4 (-0.6) 56.6 (+32.8) 42.5 (+12.2) 69.7 (+37.7) 20.8 (+6.3)
SFT + Rubric-based RL 36.5 (+28.0) 39.7 (+11.5) 57.1 (+1.8) 54.1 (-0.9) 57.4 (+33.6) 43.2 (+12.9) 71.6 (+39.6) 23.7 (+9.2)
SFT + RuscaRL (Ours) 35.4 (+26.9) 42.7 (+14.5) 58.2 (+2.9) 55.1 (+0.1) 57.7 (+33.9) 42.6 (+12.3) 72.0 (+40.0) 23.1 (+8.6)
RuscaRL consistently outperforms Rubric-based methods across tasks. As shown in Table 2, in the direct RL setting, RuscaRL achieves the best performance on most medical, writing, and instruction-following tasks, delivering larger and more stable gains than Rubric-based RL and MeRF (e.g., 56.4 vs. 52.0 and 36.8 accuracy on HealthBench-500 with Qwen2.5-7B-Instruct). Compared to exploration-enhanced methods RL-Plus and EA-RL, RuscaRL also demonstrates superior performance, indicating that our scaffolding mechanism provides more effective exploration guidance than entropy-based or hybrid-policy approaches. In the SFT-then-RL setting, both RuscaRL and Rubric-based RL achieve additional improvements over SFT, with RuscaRL yielding larger gains across most tasks, though the magnitude is smaller than in the direct RL setting. We argue that RuscaRL essentially leverages rubrics as prior knowledge to guide exploration, while SFT also serves to accelerate RL exploration (cold start). Since both mechanisms overlap in facilitating exploration, this may explain why the performance gap between RuscaRL and Rubric-based RL narrows under the SFT-then-RL setting.
@@ 5.3 Analysis
In this subsection, we conduct a detailed analysis of RuscaRL using Qwen2.5-7B-Instruct as the initial model and HealthBench as the training and evaluation dataset. In addition, we compare RuscaRL and Rubric-based RL in the following analysis. We use the Best-of-NFor cost considerations in the Best-of-N evaluation, we use Qwen3-32B (non-thinking) as the Grader LLM. metric to reflect both the model’s reasoning boundaries (at large N) and sampling efficiency (at small N).
RuscaRL substantially improves sampling efficiency and reasoning boundaries. As shown in Figure 3(a), RuscaRL substantially improves single-sample quality at N=1, indicating that the scaffolding mechanism effectively enhances the model’s reasoning stability. At N=2048, its performance ceiling surpasses both the initial model and Rubric-based RL, validating its advantage in expanding the reasoning boundary. Moreover, RuscaRL exhibits a steeper performance curve across N, meaning it can achieve the same performance with fewer samples. Further analysis (Appendix D.5.1) indicates that RuscaRL also produces highly novel responses that the initial model could barely generate, showing that rubric scaffolding effectively breaks the exploration bottleneck and discovers new solutions.
RuscaRL achieves exploration-exploitation balance. As shown in Figure 3(b), RuscaRL demonstrates a well-balanced exploration–exploitation trajectory: policy entropy first rises as the model explores diverse reasoning trajectories, then declines as it converges to high-quality patterns. In contrast, Rubric-based RL plateaus under continuous entropy decline. Validation accuracy (Figure 3(c)) consistently shows RuscaRL achieving the best performance throughout training, demonstrating sustained improvement without bottlenecks. Similar trends are observed in Self-BLEU and Semantic Distance (Appendix D.5.2), confirming that RuscaRL achieves effective exploration followed by stable exploitation.
(a) Best-of-N performance
(b) Training policy entropy
(c) Validation evaluation score
(d) Sigmoid decay function
Figure 3: Best-of-N evaluation and training dynamics. We compare RuscaRL and Rubric-based RL across sampling efficiency, policy entropy, validation score, and the sigmoid decay schedule.
@@ 5.4 Ablation Studies
Inter-step Decay Analysis. We then analyze different decay functions for inter-step control during training. We formally define the base scaffolding intensity of inter-step control as $f(t)$, where $t$ is the normalized training progress ($t\in[0,1]$). We compare the following decay functions: (1) Sigmoid (Ours): S-shaped decay function $f(t)=\frac{1}{1+e^{\alpha(t-t_{0})}}$, where parameter $\alpha$ controls the steepness of decay and $t_{0}$ controls the midpoint of decay, achieving smooth nonlinear transitions. (2) Constant: Constant control $f(t)=1$, maintaining constant full scaffolding throughout training. (3) Linear: Linear decay function f(t)=1tf(t)=1-t, achieving uniform linear decrease. (4) Power (n): Power decay function $f(t)=(1-t)^{n}$, where n controls the curvature of decay, including various power configurations.
Based on the superior performance of the sigmoid function, we further analyze the effects of both parameter dimensions (speed $\alpha$ and midpoint $t_{0}$) using Qwen2.5-7B-Instruct as the initial model and HealthBench as the training and evaluation dataset. Figures 4(c) and 4(d) demonstrate the performance differences across various sigmoid parameter configurations, ultimately determining the optimal configuration as $\alpha=125,t_{0}=0.2$. (1) Removing scaffolding too fast (large $\alpha$) prevents the model from adapting to rapid scaffolding changes, easily causing training instability; while removing scaffolding too slowly (small $\alpha$) leads to incomplete early-stage scaffolding, failing to fully stimulate the model’s exploration capability, and prolonged retention of scaffolding in later stages also causes overfitting issues. (2) Starting decay too early (small $t_{0}$) leads to insufficient scaffolding support, causing the model to lack necessary guidance in early training stages; while starting decay too late (large $t_{0}$) causes the model to over-rely on scaffolding, ultimately resulting in overfitting. Notably, RuscaRL maintains strong performance and consistently outperforms Rubric-based RL across a wide range of hyperparameters, with degradation only at extreme values.